How do you simplify (i)-5+(5+8i)? I tried distributing the i to the -5 & got -5i+0i-5. My teacher said that was wrong. Can you please help? Thanks!
Answer by tommyt3rd(5050)

(i)-5+(5+8i)=
i-5+5+8i=
9i